the sharp razor company manufactures two types of electric shavers. one shaver is cordless and requires 5 hours to make. the other shaver has an electric cord and takes 3 hours to make. the company has only 16 work hours per day to manufacture shavers. the despatch department can pack and ship at most 300 shavers per day. let x be the number of cordless shavers made per day and y be the number of shavers with a cord manufactured per day. what is the correct system of linear equations that represent this situation??
Answer by addingup(3677)

x+y= 300 therefore x= 300-y
5x+3y=16
We can substitute in the first equation:
300-y+y= 16
300= 16 and this is what the problem says, they can pack at most 300 and they work 16 hours.